OpenStudy (anonymous):

What is the exact value of the expression . 6 sqrt y^8-sqrt 9y^4+ 3 sqrt y^8 - sqrt y^4 Simplify if possible.

OpenStudy (radar):

\[6\sqrt{y ^{8}}-\sqrt{9y ^{4}}+3\sqrt{y ^{8}}-\sqrt{y ^{4}}=\]\[6y ^{4}-3y ^{2}+3y ^{4}+3y ^{4}-y ^{2}=12y ^{4}-4y ^{2}=4y ^{2}(3y ^{2}-1)\]